OpenStudy (anonymous):

How many solutions are there to the following system of equations? 4x - 8y = 0 -3x + 8y = -7 A. 0 B. 1 C. 2 D. infinitely many

OpenStudy (anonymous):

im not sure bt i say a

OpenStudy (anonymous):

put them in slope intercept form -8y=-4x y=1/2x -3x+8y=-7 8y=3x-7 y=3/8x-7/8 they have different slopes so theyll have one solution

OpenStudy (anonymous):

he seems smarter :D

OpenStudy (anonymous):